OpenStudy (anonymous):

Find two consecutive even integers whose sum is -50. Which of the following equations could be used to solve the problem? A.2x = -50 B.2x + 1 = -50 C.2x + 2 = -50 D.x^2 + 1= -50

OpenStudy (anonymous):

C.) 2x + 2 = -50

OpenStudy (anonymous):

you sure?

OpenStudy (anonymous):

if first even no = x then consecutive even no = x+2 their sum is x + (x+2) = 2x + 2 = -50

OpenStudy (anonymous):

yeah absolutely sure....

OpenStudy (anonymous):

thanks d0ll :)

OpenStudy (anonymous):

if satisfied, pls click GOOD ANSWER button (: }))